OpenStudy (anonymous):

7th grade help plz will give medal

OpenStudy (anonymous):

The base of a triangle with height h and area A is given by the formula b= 2a/h. What is the base of a triangle with area 17 cm2 and height 5 cm?

OpenStudy (anonymous):

btw im in 9th so i think i can answer

OpenStudy (anonymous):

the base is 6.8cm

OpenStudy (anonymous):

its \[2\times17\div5\]

OpenStudy (anonymous):

it comes 6.8 cm

OpenStudy (anonymous):

Tony wants to calculate his average quiz score. The formula A= 2/n gives Tony’s average score A for a total of T points on n quizzes. Tony had 47 points on 5 quizzes. What is Tony’s average score?

OpenStudy (anonymous):

I found your question and let me think for a sec

OpenStudy (anonymous):

wait what does n stand for

OpenStudy (anonymous):

The formula d = rt gives the distance traveled in time t at rate r. If a bicyclist rides for 2.3 hours and travels 41 miles, what was her average speed? A. 9.43 miles per hour B. 5.6 miles per hour C. 17.83 miles per hour D. 38.7 miles per hour

OpenStudy (anonymous):

The formula A = l • w gives the area of a rectangle. A rectangle has an area of 37.2 cm2 and a length of 12 cm What is the width of the rectangle?

OpenStudy (anonymous):

wait I think i have some of the answers

OpenStudy (anonymous):

area=length.breath

OpenStudy (anonymous):

1) 0.4 (2) C (3) 3.142

OpenStudy (anonymous):

wht???

OpenStudy (anonymous):

let me type that better

OpenStudy (anonymous):

37.7/12=3.14 is the answer

OpenStudy (anonymous):

danish help me on the first on I got the other 2

OpenStudy (anonymous):

the second one is C

OpenStudy (anonymous):

I got an 40

OpenStudy (anonymous):

no its 3.14

OpenStudy (anonymous):

ultimate which question u r asking

OpenStudy (anonymous):

The formula d = rt gives the distance d traveled in time t at rate r. A bicyclist rides at a constant rate of 22 kmph. How many kilometers will he travel in 2.5 h?

OpenStudy (anonymous):

danish do you have any idea on the answer to the first question she asked

OpenStudy (anonymous):

wait i mean the second one

OpenStudy (anonymous):

\[22\times2.5\]

OpenStudy (anonymous):

the one invloving A=2/n

OpenStudy (anonymous):

55 answer

OpenStudy (anonymous):

ultimate i just divided 47/5

OpenStudy (anonymous):

lol i didn't even think of that